# Adaptive LPD Radar Waveform Design with Generative Deep Learning

### v1

@@ Abstract
We propose a novel, learning-based method for adaptively generating low probability of detection (LPD) radar waveforms that blend into their operating environment. Our waveforms are designed to follow a distribution that is indistinguishable from the ambient radio frequency (RF) background—while still being effective at ranging and sensing. To do so, we use an unsupervised, adversarial learning framework; our generator network produces waveforms designed to confuse a critic network, which is optimized to differentiate generated waveforms from the background. To ensure our generated waveforms are still effective for sensing, we introduce and minimize an ambiguity function-based loss on the generated waveforms. We evaluate the performance of our method by comparing the single-pulse detectability of our generated waveforms with traditional LPD waveforms using a separately trained detection neural network. We find that our method can generate LPD waveforms that reduce detectability by up to 90% while simultaneously offering improved ambiguity function (sensing) characteristics. Our framework also provides a mechanism to trade-off detectability and sensing performance.
@@ II Related Work
Low Probability of Detection. Waveform design for Low Probability of Detection (LPD) and Low Probability of Intercept (LPI) has been an area of significant interest, particularly for military applications. The primary goal of these designs is to minimize the radar’s detectability while maintaining sufficient operational performance. Early LPI/LPD waveforms were characterized by low temporal, spectral, or spatial energy densities, which made them difficult to discern from background noise [19]. Techniques such as FMCW, frequency hopping, and non-traditional modulation methods like polyphase coding were common [4]. However, these methods often came with trade-offs in terms of range resolution and target identification capabilities. Noise radar emerged as an additional LPD approach that utilized random noise-like signals to reduce detectability. These waveforms have desirable ambiguity functions and are resistant to electronic countermeasures (i.e., jamming), but their random nature requires longer dwell times and more expensive processing [20].
@@
In this work, we introduce a novel application of unsupervised generative adversarial neural networks (GANs) to adaptively generate optimal LPD radar waveforms from measurements of the RF background. By using a conditional GAN, we simplify the perception-action cycle by generating waveforms directly from measurements. Unlike supervised waveform design methods, we do not train our GAN on a dataset of radar waveforms [31, 32]. Instead, we train our GAN on unlabeled data that is representative of (or taken directly from) the target environment, and then fine-tune it with a user-defined ambiguity function to produce LPD waveforms with desirable ambiguity characteristics. These LPD waveforms match the statistical behavior of the RF background distribution, rather than relying on assumptions about the environment or detector to reduce detectability (e.g., white noise, frequency hopping). This flexible framework reduces the data burden associated with deep learning-based techniques while enabling the model to adapt to changing RF environments or operational requirements.
@@ III-A Waveform Design
A good LPD waveform is therefore one that is similar to the RF background distribution and has a desirable ambiguity function. We can set up this optimization problem by minimizing the KLD between generated and background signals, and by minimizing the distance between the generated ambiguity function and some desired target ambiguity function. In this work, we accomplish this optimization with deep neural networks.
GeneratorRefer to caption
CriticRefer to caption
@@ III-B Generative Adversarial Networks
Deep generative models (DGMs) are a type of machine learning model that aim to learn the underlying probability distribution of complex and high-dimensional data [16]. They can then generate new and representative samples from the learned distribution. For example, a DGM trained on an image dataset of human faces can generate realistic faces that do not actually exist [35]. Common DGMs include variational autoencoders (VAEs) [36], generative adversarial networks (GANs) [15], normalizing flows [37], and diffusion models [35]. Although each method has strengths, we utilize GANs in this work as the adversarial training framework is particularly well suited to our LPD waveform design problem.
@@
By optimizing (10), we force the generator’s distribution $\mathbb{P}_{g}$ to match the real distribution $\mathbb{P}_{r}$. We can use this Wasserstein-1 minimization in place of KLD for LPD waveform generation by setting the real distribution equal to the RF background distribution ($\mathbb{P}_{r}=\mathbb{P}_{b}$) for a target environment. This will force our generator to learn to generate waveforms that are statistically similar—have small Wasserstein-1 distance—to the RF background.
In the context of RF background signals, it is crucial to note that the statistical properties of the distribution $\mathbb{P}_{b}$ are time-dependent. Over long timescales—i.e., minutes to hours—the background exhibits consistent, cyclical behaviors. However, over shorter timescales—i.e., microseconds to milliseconds—the background is inherently nonstationary, with rapidly fluctuating characteristics. To reconcile these disparate timescales and to adapt to the nonstationary nature of short-term RF background signals, we employ a conditional WGAN-GP (cWGAN-GP) model. By doing so, we enable the generator to produce waveforms from the conditional distribution $\mathbb{P}_{g|y}$, where $y$ is an instantaneous RF background measurement taken at or near the time of inference. This conditional approach allows us to train the WGAN on a long-term, quasi-stationary distribution while generating examples that are instantaneously conditioned on the most recent sample, thereby providing a more accurate Wasserstein-1 approximation in the nonstationary regime. This gives us the cWGAN-GP objective:
@@ 4 Combined Barker 16 [1, 8]
$*{}^{*}$ Combined Barker 169 [10.58, 84.6]
${}^{{\dagger}}$ Max. Len. Pseudo Rand. 15 [1, 4.5]
${}^{{\dagger}}$ Max. Len. Pseudo Rand. 31 [0.235, 10.5]
${}^{{\dagger}}$ Max. Len. Pseudo Rand. 63 [4.221, 18.9]
@@
${}^{{\dagger}}$ T1 N/A [2, 16]
${}^{{\dagger}}$ T2 N/A [1.5, 12]
${}^{{\dagger}}$ T3 N/A [1, 8]
${}^{{\dagger}}$ Polyphase Barker 7 [0.875, 7]
${}^{{\dagger}}$ Polyphase Barker 20 [1, 8]
${}^{{\dagger}}$ Polyphase Barker 40 [2, 16]
${}^{{\dagger}}$ P1 N/A [5, 20]
${}^{{\dagger}}$ P2 N/A [3.2, 25.6]
${}^{{\dagger}}$ P3 N/A [3.2, 25.6]
${}^{{\dagger}}$ P4 N/A [5, 20]
@@
$*{}^{*}$lass omitted from experiments ${}^{{\dagger}}$ modulation
@@
where $B$, the sweep bandwidth, is a discrete random variable drawn from a uniform distribution of integer values $B\sim U(5,19)$. We use a normalized pulse duration $T=1$, such that the time-bandwidth product $BT$ of our generated dataset lies in the range of $[5,19]$. We introduce additive white Gaussian noise as $n(t)\sim\mathcal{N}(0,\sigma^{2})$. Our generated dataset contains 50,000 waveforms with 1024 IQ time samples at 30 dB SNR. We use this simple chirp dataset to demonstrate the distribution-matching capabilities of our WGAN.
@@
Each class contains 106,496 waveforms with variable SNRs from –20 to 30 dB in increments of 2 dB. Each waveform consists of 1024 IQ time samples.
@@ V-B LPD Effectiveness
Figure 5: Evaluation Framework. Training process for the detector neural network used to evaluate LPD performance. We add two additional classes to SIDLE—generated and empty—to create the combined dataset. We add randomly time-shifted RF background signals to these examples using waveforms from RadioML 2018.01A. We train a ResNet-18 detection network to identify if a radar signal is present or not.
We evaluate the success of our LPD waveform design method by comparing the detectability of our generated waveforms to common radar waveforms from the Air Force Research Laboratory’s Signal Instantiations for Deep Learning Experiments (SIDLE) dataset [43, 45, 46]. This dataset contains simulations of 23 radar modulations, standard and LPD, with variable pulse widths shown in Table I. Each class contains 11,000 clean IQ waveforms, and we randomly select a 1024-sample segment from each waveform. For waveforms with less than 1024 samples, we randomly delay the waveform within a 1024 sample frame. We omit class 6 from this dataset as its longer pulse widths required too much memory to use effectively.
We compare the detectability of generated and SIDLE waveforms by training a separate detection neural network and evaluating its performance on the binary detection problem. Neural networks have been shown to be excellent modulation classifiers [47, 48, 49], and generating waveforms that can deceive them (without explicit adversarial attack) is a difficult task.
The detector network’s goal is to determine if any radar signal is present within a given input signal in the presence of RF background noise (i.e., it is performing detection, not classification). We train a ResNet-18 neural network [50] on a combined dataset equally composed of radar waveforms (SIDLE and generated) and an empty class with no signals present. Examples from the combined dataset are combined with randomly time-shifted RF background signals from the RadioML dataset at 0 dB SNR, and are then standardized to 0 mean, 1 standard deviation. This process is summarized in Fig. 5.
We train the detector with an 80%/20% train/test split—sampled equally across SNRs and classes—and we generate new waveforms from our generator with every training and testing epoch. This exposes the detector to the full distribution of generated waveforms $\mathbb{P}_{g}$, rather than a subset of fixed waveforms. We randomly time shift the RF background signals to ensure our generator does not have oracle time information. We also split the RadioML dataset into two equal parts: one for training our generator and a second for training the detector. This ensures our generator is not trained on any waveforms used to evaluate its LPD effectiveness.
@@ V-C Sensing Metrics
We evaluate the sensing performance of our generated waveforms by comparing their ambiguity characteristics to the SIDLE waveforms. The goal of our ambiguity loss is to optimize our generated waveforms toward the ideal “thumbtack” ambiguity. The resulting waveforms should have low ambiguity in delay (range) and Doppler (velocity) with low sidelobes. We use three ambiguity function metrics to evaluate sensing performance: mainlobe width, peak sidelobe level (PSL), and integrated Doppler ambiguity. The mainlobe width and PSL evaluate performance in the zero-Doppler regime, and the integrated Doppler evaluates the overall Doppler ambiguity response of the waveforms.
@@ VI Experimental Results
Figure 8: Detection Rate vs Data Quantity. Detection rate comparison for SIDLE T3 and generated waveforms with variable training examples seen by the detection network. Standard deviation over 10 runs shown as vertical lines. Generated waveforms significantly outperform best SIDLE waveform (T3) even with 10$\times$ more training data.
@@ VI-A Qualitative Example
Figure 9: ROC Curve. Receiver operating characteristic (ROC) curves for a supervised binary detection network tasked with detecting the SIDLE T3 waveforms and our generated waveforms. Our generated waveforms are strictly harder to detect.
Figure 10: Tunable Sensing Performance. Shown are detectability vs sensing metrics for LPD generators that have been fine-tuned with variable ambiguity loss strengths (shades of green). They are compared to the baseline SIDLE classes (orange), and a baseline generator without ambiguity loss fine-tuning (blue). Leftmost plot’s x-axis represents mainlobe width—lower implies better range resolution. Center plot’s x-axis represents peak sidelobe level—lower implies better target separation. Rightmost plot’s x-axis represents integrated Doppler ambiguity—lower implies less velocity ambiguity. Adjusting the ambiguity loss strength ($\eta$) allows one to control the trade-off between sensing ability and detectability. Increasing the ambiguity loss strength yields a less ambiguous waveform with narrower mainlobes, improved peak sidelobe level, and lower Doppler ambiguity at the cost of increased detectability. All values of ambiguity loss strength yield waveforms that offer comparable or improved sensing metrics compared to the SIDLE classes while being harder to detect.
@@ VI-B Detectability
We test this hypothesis by training our conditional WGAN-GP on the 50% training split of RadioML with the cWGAN-GP loss (Eq. 11). We omit the ambiguity loss to isolate the effects of the cWGAN-GP loss term. We compare the performance of a separately trained detection network on detecting our generated waveforms and common radar waveforms from the SIDLE dataset. These waveforms are summed to the 50% test split of RadioML at 0 dB SNR.
We find that our generator is able to produce effective LPD waveforms. The generator produces new waveforms with every training epoch, so its effectiveness is dependent on the amount of training examples seen by the detector, as shown in Fig. 8. For an equal number of training examples as the SIDLE classes ($8800$), our generated waveforms maintain an average detection rate of $20.8\%\pm 8.1\%$. All SIDLE examples are detected with nearly $100\%$ accuracy—the T3 waveforms are the most difficult to detect with a detection rate of $99.7\%$.
When the detector is exposed to 10$\times$ more generated data (88,000 training examples), detection rate is increased to $74.5\%$. This is expected, as prolonged exposure to generated waveforms will allow the detector to more accurately learn the generated distribution. However, this behavior plateaus, with 100$\times$ more generated data yielding an $80.8\%$ detection rate. This shows that the generated waveforms maintain improved LPD performance even with significantly more training data.
We compare the receiver operating characteristic (ROC) curves for the generated and T3 waveforms in Fig. 9 using a detector trained on an equal number of examples for SIDLE and generated waveforms (8800 per class). To increase the false positive rate for these curves, we retrain the detector multiple times with a decreasing ratio of empty labels in the training dataset. We find that the T3 curve reflects the behavior of a perfect detector, and the generated waveforms significantly reduce the performance of the detector.

### v2

@@ Abstract
We propose a learning-based method for adaptively generating low probability of detection (LPD) radar waveforms that blend into their operating environment. Our waveforms are designed to follow a distribution that is indistinguishable from the ambient radio frequency (RF) background—while still being effective at ranging and sensing. To do so, we use an unsupervised, adversarial learning framework; our generator network produces waveforms designed to confuse a critic network, which is optimized to differentiate generated waveforms from the background. To ensure our generated waveforms are still effective for sensing, we introduce and minimize an ambiguity function-based loss on the generated waveforms. We evaluate the performance of our method by comparing the single-pulse detectability of our generated waveforms with traditional LPD waveforms using a separately trained detection neural network. We find that our method can generate LPD waveforms that reduce detectability by up to 90% while simultaneously offering improved ambiguity function (sensing) characteristics. Our framework also provides a mechanism to trade-off detectability and sensing performance.
@@ II Related Work
Low Probability of Detection. Waveform design for Low Probability of Detection (LPD) and Low Probability of Intercept (LPI) has been an area of significant interest, particularly for military applications. The primary goal of these designs is to minimize the radar’s detectability while maintaining sufficient operational performance. Early LPI/LPD waveforms were characterized by low temporal, spectral, or spatial energy densities, which made them difficult to discern from background noise [19]. Techniques such as FMCW, frequency hopping, and non-traditional modulation methods like polyphase coding were common [4]. However, these methods often came with trade-offs in terms of effective range, resolution, or other factors such as complexity and cost [4]. Noise radar emerged as an additional LPD approach that utilized random noise-like signals to reduce detectability. These waveforms have desirable ambiguity functions and are resistant to electronic countermeasures (i.e., jamming), but their random nature requires longer dwell times and more expensive processing [20].
@@
In this work, we introduce a novel application of unsupervised generative adversarial networks (GANs) to adaptively generate optimal LPD radar waveforms from measurements of the RF background. By using a conditional GAN, we simplify the perception-action cycle by generating waveforms directly from measurements. Unlike supervised waveform design methods, we do not train our GAN on a dataset of radar waveforms [31, 13]. Instead, we train our GAN on unlabeled data that is representative of (or taken directly from) the target environment, and then fine-tune it with a user-defined ambiguity function to produce LPD waveforms with desirable ambiguity characteristics. These LPD waveforms match the statistical behavior of the RF background distribution, rather than relying on assumptions about the environment or detector to reduce detectability (e.g., white noise, frequency hopping). This flexible framework reduces the data burden associated with deep learning-based techniques while enabling the model to adapt to changing RF environments or operational requirements.
@@ III-A Waveform Design
A good LPD waveform is therefore one that is similar to the RF background distribution and has a desirable ambiguity function. We can set up this optimization problem by minimizing the distance between the distributions of generated and background signals, and by minimizing the distance between the generated ambiguity functions and some desired target ambiguity function. In this work, we accomplish this optimization with deep neural networks.
Generator
Critic
@@ III-B Generative Adversarial Networks
Deep generative models (DGMs) are a type of machine learning model that aim to learn the underlying probability distribution of complex and high-dimensional data [16]. They can then generate new and representative samples from the learned distribution. For example, a DGM trained on an image dataset of human faces can generate realistic faces that do not actually exist [34]. Common DGMs include variational autoencoders (VAEs) [35], generative adversarial networks (GANs) [15], normalizing flows [36], and diffusion models [34]. Although each method has strengths, we utilize GANs in this work as the adversarial training framework is well suited to our LPD waveform design problem—the framework is explicitly training a generator to trick a detector, which closely mirrors the LPD challenge.
@@
By optimizing (10), we force the generator’s distribution $\mathbb{P}_{g}$ to match the real distribution $\mathbb{P}_{r}$. We can use this Wasserstein-1 minimization in place of KLD for LPD waveform generation by setting the “real” distribution equal to the empirical RF background distribution ($\mathbb{P}_{r}=\mathbb{P}_{b}$) for a target environment. This will force our generator to learn to generate waveforms that are statistically similar—have small Wasserstein-1 distance—to the RF background.
In the context of RF background signals, it is crucial to note that the statistical properties of the distribution $\mathbb{P}_{b}$ are time-dependent. Take, for example, a multi-emitter urban environment. Over long timescales—i.e., minutes to hours—the background exhibits consistent, cyclical behaviors [40]. However, over shorter timescales—i.e., microseconds to milliseconds—the background is inherently nonstationary, with rapidly fluctuating characteristics as overlapping, independent signals are emitted by various sources [41]. To reconcile these disparate timescales and to adapt to the nonstationary nature of short-term RF background signals, we employ a conditional WGAN-GP (cWGAN-GP) model. By doing so, we enable the generator to produce waveforms from the conditional distribution $\mathbb{P}_{g|y}$, where $y$ is an instantaneous RF background measurement taken at or near the time of inference. This conditional approach allows us to train the WGAN on a long-term, quasi-stationary distribution while generating examples that are instantaneously conditioned on the most recent sample, thereby providing a more accurate Wasserstein-1 approximation in the nonstationary regime. This gives us the cWGAN-GP objective:
@@ 4 Combined Barker 16 [1, 8]
6 Combined Barker 169 [10.58, 84.6]
7 Max. Len. Pseudo Rand. 15 [1, 4.5]
8 Max. Len. Pseudo Rand. 31 [0.235, 10.5]
9 Max. Len. Pseudo Rand. 63 [4.221, 18.9]
@@
13 T1 N/A [2, 16]
14 T2 N/A [1.5, 12]
15 T3 N/A [1, 8]
16 Polyphase Barker 7 [0.875, 7]
17 Polyphase Barker 20 [1, 8]
18 Polyphase Barker 40 [2, 16]
19 P1 N/A [5, 20]
20 P2 N/A [3.2, 25.6]
21 P3 N/A [3.2, 25.6]
22 P4 N/A [5, 20]
@@
Class omitted from experiments    LPD modulation
@@
where $B$, the sweep bandwidth, is a discrete random variable drawn from a uniform distribution of integer values $B~{}\sim~{}U(5,19)$ Hz. We use a normalized pulse duration $T=1$ second, such that the time-bandwidth product $BT$ of our generated dataset lies in the range of $[5,19]$. We introduce additive white Gaussian noise as $n(t)\sim\mathcal{N}(0,\sigma^{2})$. Our generated dataset contains 50,000 waveforms with 1024 IQ time samples at 30 dB SNR. We use this simple chirp dataset to demonstrate the distribution-matching capabilities of our WGAN.
@@
Each class contains 106,496 waveforms with variable SNRs from –20 to 30 dB in increments of 2 dB. Each waveform consists of 1024 IQ time samples, and SNR is calculated based on the average power across all 1024 samples.
@@ V-B LPD Effectiveness
Figure 5: Evaluation Framework. Training process for the detector neural network used to evaluate LPD performance. We add two additional classes—generated and empty—to the baseline waveforms to create the combined dataset. We add randomly time-shifted RF background signals to these examples using waveforms from RadioML 2018.01A that have not been seen by our generator network. We evaluate detectability with an FFT Accumulation Method (FAM)-based cyclostationary detector and a neural network-based detector.
We evaluate the success of our LPD waveform design method by comparing the detectability of our generated waveforms to common radar waveforms from the Air Force Research Laboratory’s Signal Instantiations for Deep Learning Experiments (SIDLE) dataset [44, 46, 47]. This dataset contains simulations of 23 radar modulations, standard and LPD, with variable pulse widths shown in Table I. The waveforms are sampled at 420 MHz, with time samples per waveform varying from 99 to 10752 samples. Each class contains 11,000 clean IQ waveforms, and we randomly select a 2.4 $\mu$s (1024 time-sample) segment from each waveform. For waveforms with less than 1024 samples, we randomly delay the waveform within a 1024 sample frame. We omit class 6 from this dataset as its longer pulse widths required too much memory to use effectively. We add an additional class to this dataset by including a randomized wideband LFM chirp with variable pulse width, offset, and sweep bandwidth ($0.2f_{s}$$0.4f_{s}$).
@@
We compare the detectability of generated and baseline waveforms with two detectors: an FFT Accumulation Method-based cyclostationary detector [5, 48] and a neural network-based detector [49, 50, 51]. Both detectors are tasked with detecting the presence of a radar waveform in the presence of the RF background. Radar waveforms are energy-normalized such that they have equivalent detection ranges, and the maximum SNR of the waveforms is 0 dB. Waveforms with pulse widths longer than 2.4 $\mu$s—such as minimum-shift keying and the P-codes—will have lower power with SNRs as low as -10 dB. These longer waveforms are normalized prior to truncation, which means the detector will receive less total energy from those waveforms. Power-normalized results with fixed waveform SNR can be found in the Appendix.
Our cyclostationary detector employs the FFT Accumulation Method (FAM) [5] to detect the cyclostationary properties inherent in radar waveforms. FAM is used to compute an estimate of the spectral correlation function (SCF) which extracts the cyclic features of the signal. Because of the variable cyclostationary characteristics of the RF background, we use an adaptive thresholding detection metric that normalizes the peak SCF value by its time-averaged values. This adaptive threshold triggers a detection when the signal exhibits a higher degree of cyclostationarity than typical, rather than triggering a detection every time a fixed threshold is passed. This performed better than alternative detection metrics (e.g., max- and mean-value) on our data.
Our neural network-based detector uses a ResNet-18 architecture [52]. We train the network on a binary detection problem using a combined dataset equally composed of radar waveforms (baseline and generated) and an empty class with no signals present. Examples from the combined dataset are energy normalized and combined with a fixed-power, randomly time-shifted RF background signals from the RadioML dataset such that the maximum SNR is 0 dB. Before training the network, inputs are standardized to 0 mean, 1 standard deviation. This process is summarized in Fig. 5.
We train the ResNet-18 detector with an 80%/20% train/test split—sampled equally across SNRs and classes—and we generate new waveforms from our generator with every training and testing epoch. This exposes the detector to the full distribution of generated waveforms $\mathbb{P}_{g}$, rather than a subset of fixed waveforms. We randomly time shift the RF background signals to ensure our generator does not have oracle time information. We also split the RadioML dataset into two equal parts: one for training our generator and a second for training the detector. This ensures our generator is not trained on any waveforms used to evaluate its LPD effectiveness.
Figure 8: FAM ROC Curve. Receiver operating characteristic (ROC) curves for the FFT Accumulation Method (FAM)-based cyclostationary detector tasked with detecting the baseline waveforms and our generated waveforms in the presence of the RF background. Shown are a subset of baseline waveforms representing the best- and worst-performing. Our generated waveforms closely mirror the random detection chance. This detector fails for low false alarm rates.
@@ V-C Sensing Metrics
We evaluate the sensing performance of our generated waveforms by comparing their ambiguity characteristics to the baseline waveforms. The goal of our ambiguity loss is to optimize our generated waveforms toward the ideal “thumbtack” ambiguity. The resulting waveforms should have low ambiguity in delay (range) and Doppler (velocity) with low sidelobes. We use three ambiguity function metrics to evaluate sensing performance: mainlobe width, peak sidelobe level (PSL), and integrated Doppler ambiguity. The mainlobe width and PSL evaluate performance in the zero-Doppler regime, and the integrated Doppler evaluates the overall Doppler ambiguity response of the waveforms.
@@
Figure 9: ResNet-18 ROC Curve. Receiver operating characteristic (ROC) curves for the supervised binary detection network tasked with detecting the baseline waveforms and our generated waveforms in the presence of the RF background. Shown are a subset of baseline waveforms representing the best- and worst-performing. Our generated waveforms are strictly harder to detect.
@@ VI-B Detectability
We test this hypothesis by training our conditional WGAN-GP on the 50% training split of RadioML with the cWGAN-GP loss (Eq. 11). We omit the ambiguity loss to isolate the effects of the cWGAN-GP loss term. We compare the performance of our FAM-based cyclostationary detector and neural network-based detector on detecting our generated waveforms and the baseline waveforms. These waveforms are energy-normalized and summed to the 50% test split of RadioML. Power-normalized results at fixed waveform SNR can be found in the Appendix.
We find that our generator is able to produce effective LPD waveforms. We compare the receiver operating characteristic (ROC) curves for both the FAM and ResNet-18 detectors, shown in Fig. 8 and Fig. 9 respectively. The ResNet-18 detector is trained on an equal number of examples for baseline and generated waveforms (8800 per class). To increase the false positive rate for the ResNet-18, we retrain the detector multiple times with a decreasing ratio of empty labels in the training dataset.
We find that the cyclostationary detector performs poorly in this detection scenario, which is expected due to the complex cyclostationary RF background. Our generated waveforms are detected at a rate of random chance at high false alarm rates, and the T3 and minimum-shift keying classes are detected at below random chance. However, this detector fails at the low false alarm rates we are most interested in, with all classes degrading to below random chance.
The ResNet-18 detector performs significantly better, especially at low false alarm rates. Nearly all baseline classes are detected at a rate of over $95\%$ at a $1\%$ false alarm rate (FAR), and that rate quickly rises to over $99\%$. The exception is the minimum-shift keying class, which is detected at a rate of $70\%$ at $1\%$ FAR. Notably, our generated waveforms are detected at a rate of $19.5\%\pm 5.5\%$ at $1\%$ FAR, and they remain more difficult to detect than the baseline waveforms at all false alarm rates.
The ResNet-18 detector is supervised, and so its performance is dependent on the amount of training examples seen by the detector, as shown in Fig. 10. For an equal number of training examples as the baseline classes ($8800$), our generated waveforms maintain an average detection rate of $24.4\%\pm 5.6\%$. Of the baseline classes, the minimum-shift keying class performs the best at a rate of $74\%$—all other waveforms are detected at a rate of over $95\%$, with most over $99\%$.
When the detector is exposed to 10$\times$ more generated data (88,000 training examples), detection rate is increased to $70.0\%$. This is expected, as increased exposure to generated waveforms will allow the detector to more accurately learn the generated distribution. However, this behavior plateaus, with 100$\times$ more generated data yielding an $80.8\%$ detection rate. This shows that the generated waveforms maintain improved LPD performance even with significantly more training data.
Figure 10: Detection Rate vs Data Quantity. ResNet-18 detection rate comparison for selected baseline waveforms and our generated waveforms with variable training quantity. Generated waveforms significantly outperform the baseline waveforms even with 10$\times$ more training data.